University of Illinois Department of Economics Spring 1998 Roger Koenker

Economics 478 Topics in Semiparametric Econometrics

This page is a link to an extended bibliography concerning semiparametric methods in econometrics. I hope that we can gradually add further references to it as the semester progresses. It is available for web viewing by clicking on the ”references” link below. Alternatively, it is available as a raw file in bibtex format by clicking on the bibtex link. In the latter mode you can use it to create your own bibliographies for more specialized projects. I would like to emphasize that an efficient way to add to such bibliographies is to exploit the features of the CIS (Current Index to Statistics) web service provided by the department of statistics at the link indicated by CIS.

References

Abadie, A., J. Angrist, and G. Imbens (1997): “Instrumental Variables Estimation of Quantile Treatment Effects,” preprint. Amemiya, T. (1982): “Two Stage Least Absolute Deviations Estimators,” Econometrica, 50, 689–712, Quantile regression for simultaneous equation models. Ammann, L. P. (1993): “Robust Singular Value Decompositions: A New Approach to Projection Pursuit,” J. of Am. Stat. Assoc., 88, 505–514. Angrist, J., G. Imbens, and D. B. Rubin (1996): “Identification of Casual Effects Using Instrumental Variables,” J. of Am. Stat. Assoc., 91, 444–472. Arcones, M. (1996): “The Bahadur-Kiefer representation of L_{p} regression estimators,” Econometric Theory, 12, 257–283, Second order asymptotics for QR. Arias, O., K. Hallock, and W. Sosa (1997): “Indidividual heterogeneity in the returns to schooling: IV-QR using Twins data,” preprint. Babu, G. (1989): “Strong represention for LAD estimators in linear models,” Prob. Theory and Related Fields, 83, 547–580. Barro, R. J., and X. Sala-i-Martin (1995): Economic Growth. McGraw-Hill: New York. Barrodale, I., and F. Roberts (1974): “Solution of an overdetermined system of equations in the ℓ_{1} norm,” Communications of the ACM, 17, 319–320. Bartels, R., and A. Conn (1980): “Linearly constrained discrete ℓ_{1} problems,” Transactions of the ACM on Mathematical Software, 6, 594–608. Bassett, G., and R. Koenker (1978): “Asymptotic Theory of Least Absolute Error Regression,” J. of Am. Stat. Assoc., 73, 618–622. Bedard, K. (1996): “School Quality and the Distribution of Male Earnings in Canada,” preprint. Bosch, R. J., Y. Ye, and G. G. Woodworth (1995): “A Convergent Algorithm for Quantile Regression With Smoothing Splines,” Comp. Stat. and Data Anal., 19, 613–630. Buchinsky, M. (1994): “Changes in US Wage Structure 1963-87: An Application of Quantile Regression,” Econometrica, 62, 405–458.

(1995): “Quantile Regression, the Box-Cox Transformation Model and U.S. Wage Structure 1963-1987,” J. of Econometrics, 65, 109–154.

(1996): “Women’s ‘return to education’ in the U.S.: Exploration by Quantile Regression with Nonparametric Sample, Selection Correction,” preprint. Chakraborty, B. (1996): “On multivariate median regression,” preprint. Chakraborty, B., and P. Chaudhuri (1995): “On an adaptive transformation-retransformation estimate of multivariate,” J. of the Royal Statistical Society (B), forthcoming. Chamberlain, G. (1994): “Quantile regression censoring and the structure of wages,” in Advances in Econometrics, ed. by C. Sims. Elsevier: New York. Chaudhuri, P. (1991): “Global Nonparametric Estimation of Conditional Quantile Functions and Their Derivatives,” J. Mult. Analysis., 39, 246–269. Chaudhuri, P. (1992): “Generalized regression quantiles: Forming a useful toolkit for robust linear regression,” in Proceedings of the Second , International Conference on Statistical Data Analysis based on the , L_{1} norm and related methods, ed. by Y. Dodge. North-Holland: New York.

(1996): “On a geometric notion of quantiles for multivariate data,” J. of the American Statistical Association, 91, 862–872. Chaudhuri, P., K. Doksum, and A. Samarov (1994): “Regression Estimators Based on Conditional Quantiles,” J. of Nonparametric Stat., 3, 317–321. Chaudhuri, P., K. Doksum, and A. Samarov (1996): “Nonparametric estimation of global functionals based on quantile regression,” preprint.

(1997): “On average derivitive quantile regression,” Annals of Stat., 25, 715–744.

Chay, K. (1995): “Evaluating the impact of the 1994 Civil Rights Act on the Economic Status of Black Men Using Censored Longitudinal, Earnings Data,” preprint. Cox, D. (1972): “Regression Models with Life Tables,” J. of the Royal Statistical Society (B), 34, 187–220. Csorgo, M., and P. Revesz (1984): “Two approaches to constructing simultaneous confidence bounds for quantiles,” Prob. Math. Statist, 4, 221–236. de Jongh, P. J., T. de Wet, and A. H. Welsh (1988): “Mallows-type Bounded-influence-regression Trimmed Means,” J. of Am. Stat. Assoc., 83, 805–810. Efron, B. (1991): “Regression percentiles using asymmetric squared error loss,” Statistica Sinica, 1, 93–125. Efron, B. (1992): “Poisson Overdispersion Estimates Based on the Method of Asymmetric Maximum Likelihood,” J. of Am. Stat. Assoc., 87, 98–107. Engle, R. F. (1982): “Autoregressive conditional heteroscedasticity with estimates of the variance of United Kingdom inflation,” Econometrica, 50, 987–1007. Englund, J.-E., U. Holst, and D. Ruppert (1988): “Recursive M-estimators of Location and Scale for Dependent Sequences,” Scand. J. of Stat., 15, 147–159. Fenton, V., and A. Gallant (1996): “Qualitative and Asymptotic Performance of SNP Density Estimators,” J. Econometrics, 74, 77 –118. Fiacco, A., and G. McCormick (1968): Nonlinear Programming: Sequential Unconstrained Minimization Techniques. Wiley: New York. Fitzenberger, B. (1996): “A Guide to Censored Quantile Regressions,” in Handbook of Statistics, ed. by C. Rao, and G. Maddala. North-Holland: New York.

(1997): “Computational aspects of censorted quantile regression,” in L_{1}-Statistical Procedures and Related Topics, ed. by Y. Dodge, IMS Lecture Notes-Monograph Series. Frisch, R. (1956): “La Résolution des problèmes de programme linèaire par la mèthode du potential logarithmique,” Cahiers du Seminaire d’Econometrie, 4, 7–20. Gallant, A., and G. Tauchen (1997): “The Relative Efficiency of Method of Moments Estimators,” preprint. Geweke, J. (1989): “Exact predictive densities for linear models with ARCH disturbances,” J. of Econometrics, 40, 63–86. Gill, P., W. Murray, M. Saunders, T. Tomlin, and M. Wright (1986): “On projected Newton barrier methods for linear programming and an equivalence to Karmarker’s projective method,” Mathematical Programming, 36, 183–209. Ginther, D. (1994): “A Nonparametric analysis of the U.S. earnings distribution,” preprint. Goldberg, P. (1996): “Dealer Price Discrimination in New Car Purchases Evidence from the Consumer Expenditure Survey,” Journal of Political Economy, 104, 622–654. Gonzago, C. (1992): “Path-following methods for linear programming,” SIAM Review, 34, 167– 224. Gutenbrunner, C. (1994): “Tests for heteroscedasticity based on regression quantiles and regression rank scores,” in Asymptotic Statistics: Proceedings of the Fifth Prague Symposium, ed. by P. Mandl, and Huskova. Physica-Verlag: Heidelberg.

(1997): “Regression Rank Statistics,” in L_{1}-Statistical Procedures and Related Topics, ed. by Y. Dodge, IMS Lecture Notes-Monograph Series. Gutenbrunner, C., and J. Jureckova (1992): “Regression quantile and regression rank score process in the linear model and derived statistics,” Ann. Statist., 20, 305–330. Gutenbrunner, C., J. Jureckova, R. Koenker, and S. Portnoy (1993): “Tests of linear hypotheses based on regression rank scores,” J. of Nonparametric Statistics, 2, 307–33. Gutenbrunner, C., and J. Jurečková (1992): “Regression Rank Scores and Regression Quantiles,” Annals of Stat., 20, 305–330, Asymptotics for QR primal and dual processes.

Gutenbrunner, C., J. Jurečková, R. Koenker, and S. Portnoy (1993): “Tests of Linear Hypotheses Based on Regression Rank Scores,” J. of Nonparametric Stat., 2, 307–331, Asymptotics for QR rank tests. Hahn, J. (1995): “Bootstrapping quantile regression models,” Econometric Theory, 11, 105–121. Hajek, J., and Z. Sidak (1967): Theory of Rank Tests. Academia, Prague. Hall, P., and S. Sheather (1988): “On the distribution of a studentized quantile,” J. Royal Stat. Soc. (B), 50, 381–391. Han, A. (1987): “Nonparametric analysis of a generalized regression model,” J. Econometrics, 35, 303–318. Hansen, M., C. Kooperberg, and S. Sardy (1996): “Triogram Models,” J. of Am. Stat. Assoc., forthcoming. Hasan, M., and R. Koenker (1996): “Robust rank tests of the unit root hypothesis,” Eonometrica, forthcoming. Hastie, T., and W. Stuetzle (1989): “Principal Curves,” J. of Am. Stat. Assoc., 84, 502–516. He, X., and S. Portnoy (1997): “Asymptotics of the Deepest Line,” preprint. Heckman, J. (1996): “Randomization as an Instrumental Variable,” The Review of Economics and Statistics, 78, 336–340. Hendricks, W., and R. Koenker (1991): “Hierarchical spline models for conditional quantiles and the demand for electricity,” JASA, 87, 58–68. Hercé, M. (1996): “Asymptotic theory of LAD estimation in a unit root process with finite variance,” Econometric Theory, 12, 129–154. Holst, U. (1985): “Recursive Estimation of Quantiles,” in Contributions to Prob. and Stat. in Honour of Gunar Blom, ed. by G. Lindgren, pp. 179– 188.

(1987): “Recursive Estimation of Quantiles Using Recursive Kernel Density Estimators,” Sequential Analysis, 6, 219–237. Honoré, B., and J. Powell (1992a): “Pairwise Difference Estimators of Censored and Truncated Regression Models,” preprint.

(1992b): “Quantile regression under random censoring,” preprint.

Honoré, B. E., and J. L. Powell (1994): “Pairwise Difference Estimators of Censored and Truncated Regression Models,” J. Econometrics, 64, 241–278, U statistics models for quantile regression estimators. Horowitz, J. (1992): “A Smoothed Maximum Score Estimator for the Binary Response Model,” Econometrica, 60, 505–531.

(1996): “Bootstrap methods for median regression models,” Econometrica, forthcoming.

Huber, P. (1981): Robust Statistics. New York: Wiley. Huskova, M. (1997): “L_{1} test procedures for detection of change,” in L_{1}-Statistical Procedures and Related Topics, ed. by Y. Dodge, IMS Lecture Notes-Monograph Series. Imbens, G., and J. Angrist (1994): “Identification and Estimation of Local Average Treatment Effects,” Econometrica, 62, 467–475. Jureckova, J. (1991): “Tests of Kolmogorov-Smirnov type based on regression rank scores,” Transactions of the 11th Prague Conference on Information Theory and Statistical Decision Functions. Karmarker, N. (1984): “A new polynomial time algorithm for linear programming,” Combinatorica, 4, 373–395. Kiefer, J. (1959): “K-sample analogues of the Kolmogorov-Smirnov and Cramer-von Mises tests,” Ann. Math. Statist, 30, 420–447. King, G. (1997): A Solution to the Ecological Inference Problem. Harvard: Cambridge. Knight, K. (1989): “Limit theory for autoregressive parameters in an infinite variance random walk,” Canadian J. of Statistics, 17, 261–278.

(1997): “Asymptotics for L_{1} regression estimators under general conditions,” preprint.

Koenker, R. (1994a): “Confidence Intervals for Regression Quantiles,” in Asymptotic Statistics: Proceedings of the 5th Prague Symposium, pp. 349–359. Springer-Verlag. Koenker, R. (1994b): “Confidence intervals for regression quantiles,” in Asymptotic Statistics: Proceedings of the 5th Prague Symposium on Asymptotic Statistics, ed. by M. P., and M. Huskova. Physica-Verlag: Heidleberg.

(1996): “Rank Tests for Linear Models,” in Handbook of Statistics, ed. by C. Rao, and G. Maddala. North-Holland: New York.

(1997): “ℓ_{1} computation: an interior monologue,” in L_{1}-Statistical Procedures and Related Topics, ed. by Y. Dodge, IMS Lecture Notes-Monograph Series. Koenker, R., and G. Bassett (1978): “Regression Quantiles,” Econometrica, 46, 33–50.

(1982a): “Robust tests for heteroscedasticity based on regression quantiles,” Econometrica, 50, 43–61. Koenker, R., and G. Bassett (1982b): “Tests of linear hypotheses and l_{1} estimation,” Econometrica, 50, 1577–1584. Koenker, R., and V. d’Orey (1987): “Computing Regression Quantiles,” Applied Statistics, 36, 383–393.

(1993): “A Remark on Computing Regression Quantiles,” Applied Statistics, 36, 383–393. Koenker, R., and J. Machado (1996): “GMM inference when the number of moment conditions is large,” preprint.

(1998): “Goodness of fit and likelihood ratio processes for quantile regression,” preprint.

Koenker, R., J. Machado, C. Skeels, and A.H.Welsh (1995): “Momentary Lapses: Moment expansions and the robustness of minimum distance estimators,” Econometric Theory, pp. 172– 197. Koenker, R., J. Machado, C. Skeels, and A. Welsh. (1994): “A note on Amemiya’s form of the weighted least squares estimator,” Australian Journal of Statistics, 35, 155–174. Koenker, R., P. Ng, and S. Portnoy (1994): “Quantile Smoothing Splines,” Biometrika, 81, 673–80. Koenker, R., and B. Park (1996): “An interior point algorithm for nonlinear quantile regression,” Journal of Econometrics, 71, 265–283. Koenker, R., and S. Portnoy (1987): “L-estimation for Linear Models,” J. of Am. Stat. Assoc., 82, 851–857, Asymptotics for L-estimators based on QR. Koenker, R., and S. Portnoy (1990): “M Estimation of multivariate regressions,” J. of American Statistical Association, 85, 1060–68. Koenker, R., and F. Schorfheide (1994): “Quantile Smoothing Splines for global temperature change,” Climatic Change, 28, 395–404. Koenker, R., and Q. Zhao (1994): “L-Estimation for linear heteroscedastic models,” J. of Nonparametric Statistics, 3, 223–235.

(1996): “Conditional quantile estimation and interence for ARCH models,” Econometric Theory, 12, 793–813. Koltchinskii, V. (1997): “M-estimation, convexity, and quantiles,” Annals of Stat., 25, 435–477, Multivariate Quantiles. Koul, H., and R. Mukherjee (1994): “Regression quantiles and related processes under long range dependent errors,” J. Mult. Anal, 51, 318–337. Koul, H., and A. Saleh (1995): “Autoregression quantiles and related rank score processes,” Annals of Statistics, 23, 670–689. Lamport, L. (1994): L^{A}TEX. Addison-Wesley. Lehmann, E. (1983): Theory of Point Estimation. Wiley: New York. Liu, R. (1990): “On a notion of data depth based on random simplices,” Annals of Statistics, 36, 405–411. Lustig, I., R. Marsden, and D. Shanno (1994): “Interior point methods for linear programming: computational state of the art with discussion,” ORSA J. on Computing, 6, 1–36. Manning, W., L. Blumberg, and L. Moulton (1995): “The demand for alcohol: the differential response to price,” Journal of Health Economics, 14, 123–148. Manski, C. (1975a): “Maximum score estimation of the stochastic utility model of choice,” J. of Eonometrics, 3, 205–228.

(1985a): “Semiparametric analysis of discrete response asymptotic properties of the maximum score estimator,” J. of Eonometrics, 27, 205–228.

(1997): “Monotone Treatment Response,” Econometrica, 65, 1311–1334.

Manski, C. F. (1975b): “Maximum Score Estimation of the Stochastic Utility Model of Choice,” J. Econometrics, 3, 205–228.

(1985b): “Semiparametric Analysis of Discrete Response: Asymptotic Properties of the Maximum Score Estimator,” J. Econometrics, 27, 313–333. Manski, C. F., and T. S. Thompson (1986): “Operational Characteristics of Maximum Score Estimation,” J. Econometrics, 32, 85–108. Marden, J. (1995): Analysing and Modeling Rank Data. Chapman-Hall: New York.

Markatou, M., and J. L. Horowitz (1995): “Robust Scale Estimation in the Error-components Model Using the Empirical Characteristic Function,” Canadian J. of Stat., 23, 369–381. Marron, J. S., and M. P. Wand (1992): “Exact Mean Integrated Squared Error,” Annals of Stat., 20, 712–736. Matzkin, R. (1994): “Restrictions of economic theory in nonparametric models,” in Handbook of Econometrics, ed. by R. Engle, and D. McFadden. North-Holland: New York. Mehrotra, S. (1992): “On the implementation of a primal-dual interior point method,” SIAM J. of Optimization, 2, 575–601. Meketon, M. (1986): “Least Absolute Value Regression,” Bell Labs Technical Report. Mosteller, F., and J. Tukey (1977): Data Analysis and Regression: A Second Course in Statistics. Addison-Wesley, Reading Mass. Nevel’son, M. (1975): “On the Properties of the recursive estimates for a functional of an unknown distribution function,” in Limit Theorems of Probability Theory, ed. by P. Revesz, vol. 11 of Colloquia Mathematica Societatis Janos Bolyai. Newey, W., and J. L. Powell (1987): “Asymmetric least squares estimation and testing,” Econometrica, 55, 819–847. Newey, W. K., and J. L. Powell (1990): “Efficient Estimation of Linear and Type I Censored Regression Models Under Conditional Quantile Restrictions,” Econometric Theory, 6, 295–317, Semiparametric optimality result for QR. Niinimaa, A., and H. Oja (1995): “On the Influence Functions of Certain Bivariate Medians,” J. Royal Stat. Soc. (B), 57, 565–574. Niinimaa, A., H. Oja, and M. Tableman (1990): “The Finite-sample Breakdown Point of the Oja Bivariate Median and of the Corresponding Half-samples Version,” Stat and Prob Letters, 10, 325–328. Ninimaa, A., H. Oja, and J. Nyblom (1992): “The Oja Bivariate Median,” Applied Statistics, 41, 611–17. Oja, H. (1983): “Descriptive statistics for multivariate distributions,” Statistics and Probability Letters, 1, 327–332. Oja, H., and A. Niinimaa (1985): “Asymptotic Properties of the Generalized Median in the Case of Multivariate Normality,” J. Royal Stat. Soc. (B), 47, 372–377. Parzen, M. I., L. Wei, and Z. Ying (1994): “A resampling method based on pivotal estimating functions,” Biometrika, 81, 341–350. Perron, P. (1989): “The Great Crash, the Oil Price Shock, and the Unit-Root Hypothesis,” Econometrica, 57, 1361–1401. Phillips, P. C. B. (1991): “A Shortcut to LAD Estimator Asymptotics,” Econometric Theory, 7, 450–463, Heuristics for QR asymptotics using generalized functions. Pollard, D. (1991): “Asymptotics for Least Absolute Deviation Regression Estimators,” Econometric Theory, 7, 186–199, Asymptotics under weaker design conditions. Portnoy, S. (1984): “Tightness of the sequence of empiric cdf processes defined from regression fractiles,” in Robust and Nonlinear Time Series Analysis, ed. by J. Franke, W. Hardle, and D. Martin. Springer-Verlag: New York.

(1989): “Asymptotic behavior of the number of regression quantile breakpoints,” SIAM J. Science Statistical Computing,, 12, 867–883. Portnoy, S. (1991): “Asymptotic Behavior of Regression Quantiles in Non-stationary, Dependent Cases,” J. Mult. Analysis., 38, 100–113, Asymptotics for QR under misspecification. Portnoy, S. (1996): “Local asymptotics for quantile smoothing splines,” Annals of Statistics, forthcoming.

(1997): “On computation of regression quantiles,” in L_{1}-Statistical Procedures and Related Topics, ed. by Y. Dodge, IMS Lecture Notes-Monograph Series. Portnoy, S., and R. Koenker (1989): “Adaptive L-estimation of linear models,” Ann. Statist,, 17, 362–381. Portnoy, S., and R. Koenker (1997): “The Gaussian Hare and the Laplacian Tortoise: Computability of squared-error versus absolute-error estimators, with discusssion,” Stat. Science, 12, 279–300. Poterba, J., and K. Rueben (1995): “The Distribution of Public Sector Wage Premia: New Evidence Using Quantile Regression Methods,” NBER Working paper #4734. Powell, J. (1986a): “Censored regression quantiles,” J. Econometrics, 32, 143–155.

(1989): “Estimation of monotonic regression models under quantile restrictions,” in Nonparametric and Semiparametric Methods in Econometrics, ed. by B. W.A., J. Powell, and G. Tauchen. Cambridge U. Press: Cambridge.

(1994): “Estimation of semiparametric models,” in Handbook of Econometrics, ed. by R. Engle, and D. McFadden. North-Holland: New York. Powell, J. L. (1983): “The Asymptotic Normality of Two-stage Least Absolute Deviations Estimators,” Econometrica, 51, 1569–1576.

(1986b): “Censored Regression Quantiles,” J. Econometrics, 32, 143–155, Asymptotics for censored quantile regression. Ramsay, J., and B. Silverman (1997): Functional Data Analysis. Springer. Rocke, D. M. (1983): “Robust Statistical Analysis of Interlaboratory Studies,” Biometrika, 70, 421–431. Rousseeuw, P., and M. Hubert (1996): “Regression Depth,” preprint. Ruppert, D. (1985): “A Newton-Raphson Version of the Multivariate Robbins-Monro Procedure,” Annals of Stat., 13, 236–245. Ruppert, D., and R. Carroll (1980): “Trimmed least squares estimation in the linear model,” Journal of the American Statistical Association, 75, 828–838. Schwarz, G. (1978): “Estimating the dimension of a model,” Annals of Statistics, 6, 461–464. Serfling, R. (1980): Approximation Theorems of Mathematical Statistics. New York: Wiley. Silverman, B. (1985): “Some aspects of the spline smoothing approach to nonparametric regression curve fitting (with discussion),” JRSS(B), 47, 1–52. Small, C. (1990): “A survey of multidimensional medians,” International Statistical Review, 58, 263–277.

(1997): The Statistical Theory of Shape. Springer.

Stoker, T. (1992): Lectures on Semiparametric Econometrics. CORE Foundation: Louvain. Stone, C. (1977): “Consistent Nonparametric Regression, with discussion,” Annals of Statistics, 5, 595–645. Stone, C., M. Hansen, C. Kooperberg, and Y. Truong (1997): “Polynomial Splines and their Tensor Products in Extended Linear Modeling, with discussion,” Annals of Statistics, 25, 1371–1470. Troung, Y. (1989): “Asymptotic properties of kernel estimators based on local medians,” Annals of Statistics, 17, 606–617. Venables, W., and B. Ripley (1997): Modern Applied Statistics with Splus. Springer. Wahba, G. (1990): Spline Models for Observational Data, vol. 59. CBMS-NSF Regional Conference Series in Applied Mathematics. Wang, Y. (1994): “The L_{2} risk of the isotonic estimator,” preprint. Weiss, A. A. (1991): “Estimating Nonlinear Dynamic Models Using Least Absolute Error Estimation,” Econometric Theory, 7, 46–68, QR asymptotics for dynamic models. Welsh, A. (1987): “Kernel estimates of the sparsity function,” in Statistical Data Analysis Based on the L_{1}-norm and Related, Methods, ed. by Y. Dodge, pp. 369–377. Elsevier: New York. Welsh, A. (1993): “Robust estimation of smooth regression and spread functions and their derivatives,” preprint. Wolak, F. A. (1987): “An exact test for multiple inequality and equality constraints in the linear regression model,” Journal of the American Statistical Association, 82, 782–793. Wright, M. (1992): “Interior Methods for constrained optimization,” Acta Numerica, pp. 341– 407. Ying, Z., S. H. Jung, and L. J. Wei (1995): “Survival Analysis With Median Regression Models,” J. of Am. Stat. Assoc., 90, 178–184. Zhou, K., and S. Portnoy (1996): “Direct use of regression quantiles to construct confidence sets in linear models,” Annals of Statistics, 24, 287–306. Zwanzig, S. (1997): “On L_{1} norm estimators in nonlinear regression and in nonlinear errors in variables models,” in L_{1}-Statistical Procedures and Related Topics, ed. by Y. Dodge, IMS Lecture Notes-Monograph Series.